Two 12-V batteries with internal resistance of 0.20 ohms and 0.25 ohms respectively are joined in parallel and a resistor of 1.0 ohms resistor is placed across the terminals. Find the current ?
1
Expert's answer
2021-10-27T07:01:47-0400
"I=\frac{12}{1+\frac{0.2(0.25)}{0.2+0.25}}=10.8\ A"